iosangel:

y and x have a proportional relationship, and y = 15 when x = 3. What is the value of x when y = 4?

Mehek:

To get from x to y: x = 3 ; y = 15 You just multiply 3 * 5 right? So now if y = 4, to get that, it would be 5 * x 4 = 5x Divide to get x

iosangel:

so the answer is 5x?

Mehek:

No you need to isolate x So you need to get rid of the 5 by x To do that you just divide 4/5 = ?

loveboy:

y and x are proportional to each other, and y varies with x ,so we can say that y=kx and when x=3 ,y is 15 so, 15=3k where we get k=5 now,we have to find the value of x when y is already given. using the same concept and we have k as well so, 4=5x x=4/5=0.8

loveboy:

hope it helps :)